# Random Lévy Looptrees and Lévy Maps

### v2

@@ 5.1.   Convergence of looptrees and labels
$C^{(\lambda)}_{t}+\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}=X^{(\lambda)}_{t-}-\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}>\delta\}}\enskip\mathop{\longrightarrow}^{(d)}_{\lambda\to\infty}\enskip X_{t-}-\sum_{s\prec t}R^{t}_{s}\mathbf{1}_{\{\Delta X_{s}>\delta\}}=C_{t}+\sum_{s\prec t}R^{t}_{s}\mathbf{1}_{\{\Delta X_{s}\leq\delta\}},$
and the right-hand side further converges to $C_{t}$ as $\delta\to 0$. We shall compare the left-hand side of this display with (twice) that of (5.1). First note that as a consequence of this last convergence, if for $\varepsilon>0$ we let
$A_{\lambda,\delta,\varepsilon}=\Bigl{\{}\Bigl{|}C^{(\lambda)}_{t}+\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\Bigr{|}<\varepsilon^{-1/2}\Bigr{\}},$
then the probability of this event goes to $1$ as first $\lambda\to\infty$, then $\delta\to 0$, and finally $\varepsilon\to 0$. Unless specified, we shall below consider limits in this sense.
Let us introduce a second event:
$B_{\lambda,\delta,\varepsilon}=\Bigl{\{}\sum_{s\prec t}\Delta X^{(\lambda)}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\leq\varepsilon\Bigr{\}}.$
Recall that $\min(R^{(\lambda),t}_{s},\Delta X^{(\lambda)}_{s}-R^{(\lambda),t}_{s})\leq R^{(\lambda),t}_{s}\leq\Delta X^{(\lambda)}_{s}$. Fix a bounded and $K$-Lipschitz function $F$ for some $K>0$, then we have:
$\mathop{}\!\mathbb{E}\Bigl{[}\Bigl{|}F\Bigl{(}\frac{1}{2}C^{(\lambda)}_{t}+\sum_{s\prec t}\min(R^{(\lambda),t}_{s},\Delta X^{(\lambda)}_{s}-R^{(\lambda),t}_{s})\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\Bigr{)}-F\Bigl{(}\frac{1}{2}\Bigl{(}C^{(\lambda)}_{t}+\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\Bigr{)}\Bigr{)}\Bigr{|}\mathbf{1}_{\{B_{\lambda,\delta,\varepsilon}\}}\Bigr{]}$
$\leq\mathop{}\!\mathbb{E}\Bigl{[}K\Bigl{(}\sum_{s\prec t}\min(R^{(\lambda),t}_{s},\Delta X^{(\lambda)}_{s}-R^{(\lambda),t}_{s})\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}+\frac{1}{2}\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\Bigr{)}\mathbf{1}_{\{B_{\lambda,\delta,\varepsilon}\}}\Bigr{]}$
$\leq K\cdot\frac{3\varepsilon}{2}.$
@@
On the complementary event $B_{\lambda,\delta,\varepsilon}^{c}$, we may adapt the argument in the proof of Theorem 1.2 in [CK14]. Precisely, we can write $R^{(\lambda),t}_{s}=\Delta X^{(\lambda)}_{s}\times u^{t}_{s}$, where $\Delta X^{(\lambda)}_{s}$ is the size of the jump of $X^{(\lambda)}$ at time $s$, and in the unconditioned process, the random variables $(u^{t}_{s};s\prec t)$ are i.i.d. with the uniform law on $[0,1]$ and are independent from $(\Delta X^{(\lambda)}_{s};s\prec t)$. Consequently,
@@
$\mathrm{Var}\biggl{(}\frac{\sum_{s\prec t}\Delta X^{(\lambda)}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}u^{t}_{s}}{\sum_{s\prec t}\Delta X^{(\lambda)}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}}\bigm{|}\Delta X^{(\lambda)}_{s};s\prec t\biggr{)}=\frac{\sum_{s\prec t}(\Delta X^{(\lambda)}_{s})^{2}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}}{12(\sum_{s\prec t}\Delta X^{(\lambda)}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}})^{2}}\leq\frac{\sup_{s\prec t}\Delta X^{(\lambda)}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}}{12\sum_{s\prec t}\Delta X^{(\lambda)}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}},$
which is bounded by $\delta/(12\varepsilon)$ on the event $B_{\lambda,\delta,\varepsilon}^{c}$. The same holds if we replace $u^{t}_{s}$ by $\min(u^{t}_{s},1-u^{t}_{s})$, in which case the numerator on left represents the contribution to the looptree distance of the small jumps, the only difference is that the expectation is now $1/4$. We infer that if we let
@@
$C_{\lambda,\delta,\varepsilon}=\biggl{\{}\frac{\frac{1}{2}C^{(\lambda)}_{t}+\sum_{s\prec t}\min(R^{(\lambda),t}_{s},\Delta X^{(\lambda)}_{s}-R^{(\lambda),t}_{s})\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}}{C^{(\lambda)}_{t}+\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}}\in\Bigl{[}\frac{1/4-\varepsilon}{1/2+\varepsilon},\frac{1/4+\varepsilon}{1/2-\varepsilon}\Bigr{]}\biggr{\}},$
then
$\mathop{}\!\mathbb{P}(B_{\lambda,\delta,\varepsilon}^{c}\cap C_{\lambda,\delta,\varepsilon}^{c})\leq\frac{\delta}{L\,\varepsilon^{3}},$
for some universal constant $L$ coming from the variances. Let $g(\varepsilon)=\max(\frac{1}{2}-\frac{1/4-\varepsilon}{1/2+\varepsilon},\frac{1/4+\varepsilon}{1/2-\varepsilon}-\frac{1}{2})$. Recall the event $A_{\lambda,\delta,\varepsilon}=\{|C^{(\lambda)}_{t}+\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}|<\varepsilon^{-1/2}\}$ and recall that $F$ is $K$-Lipschitz. Then we have:
$\mathop{}\!\mathbb{E}\Bigl{[}\Bigl{|}F\Bigl{(}\frac{1}{2}C^{(\lambda)}_{t}+\sum_{s\prec t}\min(R^{(\lambda),t}_{s},\Delta X^{(\lambda)}_{s}-R^{(\lambda),t}_{s})\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\Bigr{)}-F\Bigl{(}\frac{1}{2}\Bigl{(}C^{(\lambda)}_{t}+\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\Bigr{)}\Bigr{)}\Bigr{|}\mathbf{1}_{\{A_{\lambda,\delta,\varepsilon}\cap C_{\lambda,\delta,\varepsilon}\}}\Bigr{]}$
$\leq\mathop{}\!\mathbb{E}\Bigl{[}K\cdot\Bigl{|}\frac{1}{2}C^{(\lambda)}_{t}+\sum_{s\prec t}\min(R^{(\lambda),t}_{s},\Delta X^{(\lambda)}_{s}-R^{(\lambda),t}_{s})\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}-\frac{1}{2}\Bigl{(}C^{(\lambda)}_{t}+\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\Bigr{)}\Bigr{|}\mathbf{1}_{\{A_{\lambda,\delta,\varepsilon}\cap C_{\lambda,\delta,\varepsilon}\}}\Bigr{]}$
$\leq\mathop{}\!\mathbb{E}\Bigl{[}K\cdot g(\varepsilon)\cdot\Bigl{|}C^{(\lambda)}_{t}+\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\Bigr{|}\mathbf{1}_{\{A_{\lambda,\delta,\varepsilon}\cap C_{\lambda,\delta,\varepsilon}\}}\Bigr{]}$
$\leq K\cdot g(\varepsilon)/\sqrt{\varepsilon},$
which converges to 00 as $\varepsilon\to 0$ since $g$ is approximately linear near 00.
@@
Recall that $F$ is $K$-Lipschitz and bounded, say $|F|\leq K^{\prime}$. Then by partitioning the space into the four events: $B_{\lambda,\delta,\varepsilon}\cup\{A_{\lambda,\delta,\varepsilon}^{c}\cap B_{\lambda,\delta,\varepsilon}^{c}\}\cup\{A_{\lambda,\delta,\varepsilon}\cap B_{\lambda,\delta,\varepsilon}^{c}\cap C_{\lambda,\delta,\varepsilon}\}\cup\{A_{\lambda,\delta,\varepsilon}\cap B_{\lambda,\delta,\varepsilon}^{c}\cap C_{\lambda,\delta,\varepsilon}^{c}\}$, we conclude that
$\Bigl{|}\mathop{}\!\mathbb{E}\Bigl{[}F\Bigl{(}\frac{1}{2}C^{(\lambda)}_{t}+\sum_{s\prec t}\min(R^{(\lambda),t}_{s},\Delta X^{(\lambda)}_{s}-R^{(\lambda),t}_{s})\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\Bigr{)}\Bigr{]}-\mathop{}\!\mathbb{E}\Bigl{[}F\Bigl{(}\frac{1}{2}\Bigl{(}C^{(\lambda)}_{t}+\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\Bigr{)}\Bigr{)}\Bigr{]}\Bigr{|}$
is upper bounded by:
$K\frac{3\varepsilon}{2}+2K^{\prime}\mathop{}\!\mathbb{P}(A_{\lambda,\delta,\varepsilon}^{c})+K\frac{g(\varepsilon)}{\sqrt{\varepsilon}}+2K^{\prime}\frac{\delta}{L\,\varepsilon^{3}},$
which goes to 00 as first $\lambda\to\infty$, then $\delta\to 0$, and finally $\varepsilon\to 0$. Recall from the display below (5.1) that the second expectation above tends to $\mathop{}\!\mathbb{E}[F(C_{t})]$ along these limits, hence so does the first one. This proves (5.1) by the Portmanteau theorem.
@@
Step 2: distance from an ancestor to $t$. So far, we have proved the convergence of the looptree distance from 00 to any given $t>0$. Let $s\in(0,t)$ and suppose first that $s\prec t$ in the looptree coded by $X$, namely that $X_{s-}\leq X_{r}$ for all $r\in[s,t]$. There are two cases according as wether $X$ jumps at time $s$ or not.
@@
If $\Delta X_{s}>0$, then recall from the proof of Lemma 3.3 that $X$ cannot make a local minimum at the height $X_{s-}$. In particular, either $X_{t}>X_{s-}$ or $t=\inf\{r>s\colon X_{r}=X_{s-}\}$. In the second case, this means that we can find $t^{\prime}<t$ arbitrarily close to $t$, with $X_{t^{\prime}}>X_{t}=X_{s-}$ and $X_{t^{\prime}}$ is arbitrarily close to $X_{t}$; this implies that the looptree distance between $t$ and $t^{\prime}$ is arbitrarily small and $t^{\prime}$ falls into the first case. Therefore we may assume when $\Delta X_{s}>0$ that $X_{t}>X_{s-}$ and actually $X_{r}>X_{s-}$ for all $r\in[s,t]$. In this case, as before, there are times $s^{(\lambda)}\prec t^{(\lambda)}$ for $X^{(\lambda)}$, such that $X^{(\lambda)}$ jumps at time $s^{(\lambda)}$, and such that $s$, $t$, $\Delta X_{s}$ and $R^{t}_{s}$ are the limits of the analogous quantities in $X^{(\lambda)}$ and the previous argument extends readily to prove the convergence of the looptree distance from $s^{(\lambda)}$ to $t^{(\lambda)}$ for $X^{(\lambda)}$ to that from $s$ to $t$ for $X$.
@@
Proof of uniqueness of the limits for bridges and excursions.
From the two previous proofs, we know that the looptree distances and labels associated with the bridges are tight and those associated with the unconditioned processes converge. Fix $\varepsilon\in(0,1)$ and notice that the looptree distance between all pairs $s,t\in[0,1-\varepsilon]$ is a measurable function of the trajectory of the process restricted to the time interval $[0,1-\varepsilon]$. The same holds for the covariance function of the labels on the time interval $[0,1-\varepsilon]$. Then the convergence of these looptree distances and labels for the bridge follows from the absolute continuity relation as in the proof of Lemma 5.1. This characterises the subsequential limits and we conclude from tightness.

### v3

@@ 6.1.   Convergence of looptrees and labels
$\displaystyle\enskip\mathop{\longrightarrow}^{(d)}_{\lambda\to\infty}\enskip$ $X_{t-}-\sum_{s\prec t}R^{t}_{s}\mathbf{1}_{\{\Delta X_{s}>\delta\}}=C_{t}+\sum_{s\prec t}R^{t}_{s}\mathbf{1}_{\{\Delta X_{s}\leq\delta\}},$
and the limit further converges in probability to $C_{t}$ as $\delta\to 0$. Hence, to conclude about the convergence of the looptree distance from 00 to $t$, it remains to compare this distance true distance, in $X^{(\lambda)}$, with the quantity we obtain when we replace the term $\sum_{s\prec t}\min(R^{(\lambda),t}_{s},\Delta X^{(\lambda)}_{s}-R^{(\lambda),t}_{s})\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}$ by $\frac{1}{2}\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}$.
Precisely, let us fix a bounded Lipschitz function $F$, with $\sup_{x}|F(x)|\leq K$ and $\sup_{x\neq y}|F(x)-F(y)|/(|x-y|)\leq L$ for some positive constants $K$ and $L$, and define
$D_{F}(\lambda,\delta)$
$=\Bigl{|}F\Bigl{(}\frac{1}{2}\Bigl{(}C^{(\lambda)}_{t}+\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\Bigr{)}+\sum_{s\prec t}\min(R^{(\lambda),t}_{s},\Delta X^{(\lambda)}_{s}-R^{(\lambda),t}_{s})\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}>\delta\}}\Bigr{)}$
$\quad-F\Bigl{(}\frac{1}{2}C^{(\lambda)}_{t}+\sum_{s\prec t}\min(R^{(\lambda),t}_{s},\Delta X^{(\lambda)}_{s}-R^{(\lambda),t}_{s})\Bigr{)}\Bigr{|}.$
Let us prove that $\mathop{}\!\mathbb{E}[D_{F}(\lambda,\delta)]\to 0$ as first $\lambda\to\infty$ and then $\delta\to 0$, then our claim follows from the Portmanteau theorem.
@@
Let us adapt the argument in the proof of Theorem 1.2 in [CK14]. Precisely, we can write $R^{(\lambda),t}_{s}=\Delta X^{(\lambda)}_{s}\times u^{t}_{s}$, where in the unconditioned process, the random variables $(u^{t}_{s};s\prec t)$ are i.i.d. with the uniform law on $[0,1]$ and are independent from $(\Delta X^{(\lambda)}_{s};s\prec t)$. Consequently,
@@
$\mathrm{Var}\biggl{(}\frac{\sum_{s\prec t}\Delta X^{(\lambda)}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}u^{t}_{s}}{\sum_{s\prec t}\Delta X^{(\lambda)}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}}\bigm{|}\Delta X^{(\lambda)}_{s};s\prec t\biggr{)}$ $=\frac{\sum_{s\prec t}(\Delta X^{(\lambda)}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}})^{2}}{12(\sum_{s\prec t}\Delta X^{(\lambda)}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}})^{2}}$
$\leq\frac{\delta}{12\sum_{s\prec t}\Delta X^{(\lambda)}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}},$
after bounding the sum of the squares by the sum without square times the largest element. The same holds if we replace $u^{t}_{s}$ by $\min(u^{t}_{s},1-u^{t}_{s})$, except that the expectation is now $1/4$ and the variance is $1/48$.
Fix $\lambda,\delta>0$ and $\varepsilon\in(0,1/4)$ and consider the following two events. First
$A_{\lambda,\delta,\varepsilon}=\Bigl{\{}\sum_{s\prec t}\Delta X^{(\lambda)}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\leq\varepsilon\Bigr{\}}$
controls the denominator above, so on the complement event, the variances are bounded above by $\delta/(12\varepsilon)$ and $\delta/(48\varepsilon)$ respectively. Second,
@@
$B_{\lambda,\delta,\varepsilon}=\biggl{\{}\biggl{|}\frac{\sum_{s\prec t}\min(R^{(\lambda),t}_{s},\Delta X^{(\lambda)}_{s}-R^{(\lambda),t}_{s})\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}}{\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}}-\frac{1}{2}\biggr{|}\leq\frac{3\varepsilon}{1-2\varepsilon}\biggr{\}},$
controls the ratios. It is straightforward to check that if $|a-1/4|\leq\varepsilon$ and $|b-1/2|\leq\varepsilon$, then $|a/b-1/2|\leq\max\{(1/4+\varepsilon)/(1/2-\varepsilon)-1/2,1/2-(1/4-\varepsilon)/(1/2+\varepsilon)\}=3\varepsilon/(1-2\varepsilon)$. A union bound and the Chebychev inequality then easily show that, for some universal constant $V$ (for example $5/24$) coming from the variances, it holds:
$\mathop{}\!\mathbb{P}(A_{\lambda,\delta,\varepsilon}^{c}\cap B_{\lambda,\delta,\varepsilon}^{c})\leq\frac{\delta V}{\varepsilon^{3}}.$
Let us introduce a third and last event:
$C_{\lambda,\delta,\varepsilon}=\Bigl{\{}\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}<\varepsilon^{-1/2}\Bigr{\}},$
whose probability tends to $1$ as first $\lambda\to\infty$, then $\delta\to 0$, and finally $\varepsilon\to 0$ since this sum is upper bounded by $X^{(\lambda)}_{t-}$ which converges in distribution to $X_{t-}$ which is finite.
@@
Recall the quantity $D_{F}(\lambda,\delta)$ and that we aim at showing that its expectation converges to 00 as first $\lambda\to\infty$, then $\delta\to 0$. We shall partition the space into the four events: $A_{\lambda,\delta,\varepsilon}\cup\{A_{\lambda,\delta,\varepsilon}^{c}\cap B_{\lambda,\delta,\varepsilon}\cap C_{\lambda,\delta,\varepsilon}\}\cup\{A_{\lambda,\delta,\varepsilon}^{c}\cap B_{\lambda,\delta,\varepsilon}^{c}\cap C_{\lambda,\delta,\varepsilon}\}\cup\{A_{\lambda,\delta,\varepsilon}^{c}\cap C_{\lambda,\delta,\varepsilon}^{c}\}$. Indeed, first, since $F$ is $L$-Lipschitz, then
$D_{F}(\lambda,\delta)\leq L\Bigl{|}\sum_{s\prec t}\min(R^{(\lambda),t}_{s},\Delta X^{(\lambda)}_{s}-R^{(\lambda),t}_{s})\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}-\frac{1}{2}\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}}\Bigr{|}.$
Using the triangle inequality, this is bounded above by $23L\varepsilon/$ 2 on the event $A_{\lambda,\delta,\varepsilon}$. If instead we factorise by the sum of the $R$’s, then on the event $B_{\lambda,\delta,\varepsilon}$, we have
$D_{F}(\lambda,\delta)\leq\frac{3L\varepsilon}{1-2\varepsilon}\sum_{s\prec t}R^{(\lambda),t}_{s}\mathbf{1}_{\{\Delta X^{(\lambda)}_{s}\leq\delta\}},$
which is now bounded above by $3L\sqrt{\varepsilon}/(1-2\varepsilon)$ on the event $C_{\lambda,\delta,\varepsilon}$. For the last two cases, we simply use that $F$ is bounded by $K$, so $D_{F}(\lambda,\delta)\leq 2K$ is uniformly bounded. Therefore, we have:
$\mathop{}\!\mathbb{E}[D_{F}(\lambda,\delta)]\leq\frac{3L\varepsilon}{2}+\frac{3L\sqrt{\varepsilon}}{1-2\varepsilon}+2K\mathop{}\!\mathbb{P}(A_{\lambda,\delta,\varepsilon}^{c}\cap B_{\lambda,\delta,\varepsilon}^{c})+2K\mathop{}\!\mathbb{P}(C_{\lambda,\delta,\varepsilon}^{c}).$
We know that the first probability on the right is bounded above by some universal constant times $\delta/\varepsilon^{3}$ and that the second probability converges to 00 as first $\lambda\to\infty$, then $\delta\to 0$, and finally $\varepsilon\to 0$. The same is then true for $\mathop{}\!\mathbb{E}[D_{F}(\lambda,\delta)]$ and the proof is complete.
@@
Step 2: distance from an ancestor to $t$. So far, we have proved the convergence of the looptree distance from 00 to any given $t>0$. Let $s\in(0,t)$ and suppose first that $s\prec t$ in the looptree coded by $X$, namely that $X_{s-}\leq X_{r}$ for all $r\in[s,t]$. There are two cases according as whether or not $X$ jumps at time $s$.
@@
If $\Delta X_{s}>0$, then recall from the proof of Lemma 4.3 that $X$ cannot make a local minimum at the height $X_{s-}$. In particular, either $X_{t}>X_{s-}$ or $t=\inf\{r>s\colon X_{r}=X_{s-}\}$. In the second case, this means that we can find $t^{\prime}<t$ arbitrarily close to $t$, with $X_{t^{\prime}}>X_{t}=X_{s-}$ and $X_{t^{\prime}}$ is arbitrarily close to $X_{t}$; this implies that the looptree distance between $t$ and $t^{\prime}$ is arbitrarily small and $t^{\prime}$ falls into the first case. Therefore we may assume when $\Delta X_{s}>0$ that $X_{t}>X_{s-}$ and actually $X_{r}>X_{s-}$ for all $r\in[s,t]$. In this case, as in the previous step, there are times $s^{(\lambda)}\prec t^{(\lambda)}$ such that $X^{(\lambda)}$ jumps at time $s^{(\lambda)}$, and such that $s$, $t$, $\Delta X_{s}$ and $R^{t}_{s}$ are the limits of the analogous quantities in $X^{(\lambda)}$ and the previous argument extends readily to prove the convergence of the looptree distance from $s^{(\lambda)}$ to $t^{(\lambda)}$ for $X^{(\lambda)}$ to that from $s$ to $t$ for $X$.
@@
Proof of convergence for bridges and excursions.
We know from the two previous proofs that the looptree distances and labels associated with the bridges are tight and those associated with the unconditioned processes converge. Fix $\varepsilon\in(0,1)$ and notice that the looptree distance between all pairs $s,t\in[0,1-\varepsilon]$ is a measurable function of the trajectory of the process restricted to the time interval $[0,1-\varepsilon]$. The same holds for the covariance function of the labels on the time interval $[0,1-\varepsilon]$. Then the convergence of these looptree distances and labels for the bridge follows from the absolute continuity relation as in the proof of Lemma 6.1. This characterises the subsequential limits and we conclude from tightness.